OpenStudy (anonymous):

The results of a class election are shown in the table below. There were a total of 600 votes. For each candidate write a fraction that shows number of votes/ total number of votes Jeremy 152 Bina 104 Jun 136 Gregory 208 Please don't use a calculator! Explain what you did please!

OpenStudy (anonymous):

HELPP Fast!!!

OpenStudy (anonymous):

Jeremy = 152/600 both r divisible by 4 so we get 38/150 = 19/75

OpenStudy (anonymous):

Also can you write a decimal because I forgot to write that lol thats what I meant not to use a calculator on :P

OpenStudy (anonymous):

Bina = 104/600 both r divisible by 4 so we get 26/150 = 13/75

OpenStudy (anonymous):

u want answer in decimals???

OpenStudy (anonymous):

yea change the fractions to decimals

OpenStudy (anonymous):

152 Jeremy =------- Now divide 152 by 6 and then shift the decimal left by two places 6 * 100 25.333..... = --------- = 0.253 approx. 100

OpenStudy (anonymous):

can you do the other ones too?

OpenStudy (anonymous):

104 17.333... Bina = ------ = --------- = 0.173 approx. 6*100 100 u can do the others the same way....

OpenStudy (anonymous):

Would Jun's be 0.266?

OpenStudy (anonymous):

136 Jun = ------ 6*100 208 Gregory = ------ 6*100 follow the same procedure as for the first two....

OpenStudy (anonymous):

Jun will be 0.2266666..... so it will be 0.227 approx

OpenStudy (anonymous):

Gregory will be 0.3466666..... so it will b 0.347 approx

OpenStudy (anonymous):

k thanks